find the equation of circle passing through the points (1,1),(2,1)and (3,2)

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

DEAR RAHUL,

The standard form of a circle whose centre is (g,f) and radius r is:
x² + y² -2gx -2gy + c = 0, where r² = g² + f² - c

First, 'plug in' the co-ords. of the first given point:
(1)² + (1)² -2g(1) - 2f(1) + c = 0
  1+ 1 -2g -2f + c =0
2-2g-2f+c=0

Repeat for the next 2 given points. You obtain 3 simultaneous linear equations for g,f,c.
Solve them and 'plug' the values for g, f and c back into the standard equation.
Hence, the eq'n. of the circle is thus obtained.
{Incidentally, so is the radius: √[g² + f² -c] }

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

sagarsingh24.iitd@gmail.com